OpenStudy (anonymous):

The equations of two lines are x - 3y = 6 and y = 3x + 2. Determine if the lines are parallel, perpendicular or neither. i got parallel as an answer?

OpenStudy (anonymous):

put equations of two line @ the same form: y=3x+2 y=1/3 x -2 (you re-wrote first one in this form; see how?) the slope = coefficient in a front of x the slope of the first line m1=3 the slope of the second line m2=1/2 parallel lines have the same slope -> your lines are not; perpendicular lines: m1=-1/m2; NOT perpendicular ; the answer: neither